# SAMSA: Efficient Transformer for Many Data Modalities

### v1

@@ 4.2. Differentiable Top-$k$ Sampling
Gumbel-Softmax (Jang et al., 2017) provides a differentiable stochastic choosing function via continuous relaxation of hardmax function. The sampling-with-replacement procedure via Gumbel-Softmax with reparameterization trick is given as follows:
$z$ $=\operatorname{ImportanceScore}(\mathbb{X}),$
$s(\mathbf{X},g)$ $=\mathbf{X}_{\operatorname{argmax}(z+g)},$
$\nabla s(\mathbf{X},g)$ $=\nabla\left(\sum_{i=1}^{n}\frac{\exp\left(\frac{z_{i}+g_{i}}{\tau}\right)}{\sum_{j=1}^{n}\exp\left(\frac{z_{j}+g_{j}}{\tau}\right)}\mathbf{X}_{i}\right),$
$S(\mathbf{X},k)$ $=\operatorname{Stack}_{i=1}^{k}(s(\mathbf{X},g)),$
where $S$, $s$, $\mathbf{X}$, $z$, $\tau$, $k$, and $g$ are the sampling function, the sampling one-vector function, the value matrix, the importance score, the Gumbel temperature hyperparameter, the number of sampled vectors, and Gumbel random variable, respectively. Since the selected vectors are independent, parallelization is trivial in sampling with replacement. This is the most important characteristic as our intended use-case of differentiable sampler in this work is as a component of deep transformer networks.
@@
While parallelization of sampling with replacement is so much desirable, the existence of duplicates harms expressivity. Therefore, we constructed a more desirable sampling method via reparameterization that rules out duplicates, is differentiable, and most importantly has great parallelizability. Our method estimates the gradient of hard sampling via a proxy soft-sampler. This soft-sampler selects top-2k important vectors. Then, for each of the top-k most important vectors, it compares against every of the k-less important vectors in k pairs with Gumbel-Sigmoid to generate gradients (a variant of Gumbel-Softmax where the number of vectors is two). While the top-k procedure is non-differentiable, putting vector pairs in probability bins makes those vectors compete for slots to be selected. As a result, through gradient descent, the more important vectors defined via data will have their score go toward up while the others go down. Mathematically, the method is described as follows:
$z$ $=\operatorname{ImportanceScore}(\mathbb{X})+g,$
$i$ $=\operatorname{ArgTopk}(z,2k),$
$p$ $=\operatorname{Sigmoid}\left(\frac{z_{i[k:]}\mathbb{1}_{k}^{\top}-\mathbb{1}_{k}z_{i[k:]}^{\top}}{\tau}\right),$
$S(\mathbf{X},k)$ $=\mathbf{X}_{i[:k]},$
$\nabla S(\mathbf{X},z,g)[j]$ $=\nabla\left(\sum_{g=1}^{k}\frac{\mathbb{X}[i[j]]*p[j,g]+\mathbb{X}[i[k]]*(1-p[j,g])}{k}\right)$
where $\mathbb{1}_{k}$ are the all-one vectors of length $k$.
Algorithm 2 Differentiable Sampling Without Replacement
0:  Importance score $\mathbb{S}\in\mathbb{R}^{n}$, number of sampled token $k$
0:  Sampling matrix $\mathbb{P}\in\mathbb{R}^{k\times n}$
1:  Select top-2k important score
2:  $\mathbb{I}\leftarrow\operatorname{ArgTopK}(\mathbf{S},2k)$
3:  Compute the proxy function
4:  $\mathbb{P}\leftarrow\mathbb{0}_{k\times n}$
5:  $\mathbb{S}\leftarrow\exp{(\mathbb{S})}$
6:  $\mathcal{S}\leftarrow\mathbb{I}_{:k}$
7:  $\mathcal{R}\leftarrow\mathbb{I}_{(k+1):2k}$
8:  $\sigma_{\mathcal{R}}\leftarrow 0$
9:  for $r$ in $\mathcal{R}$ do
10:     $\sigma_{\mathcal{R}}\leftarrow\sigma_{\mathcal{R}}+\mathbb{S}_{r}$
11:  end for
12:  for $i$ from $1$ to $k$ do
13:     $s\leftarrow\mathcal{S}_{i}$
14:     $\mathbb{P}_{i,s}\leftarrow\mathbb{S}_{s}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
15:     for $r$ in $\mathcal{R}$ do
16:        $\mathbb{P}_{i,r}\leftarrow\mathbb{S}_{r}/(\mathbb{S}_{s}+\sigma_{\mathcal{R}})$
17:     end for
18:  end for
19:  Return $(\operatorname{Onehot}(\mathbb{I}_{:k})-\mathbb{P}).\operatorname{detach()}-\mathbb{P}$
4.3. Integration to Self-Attention Mechanism
@@
In the previous section, we introduced our differentiable top-k sampling method, designed to enhance the efficiency of the self-attention mechanism. A straightforward integration of this method would involve allowing all query vectors to attend to a subset of sampled key and value vectors. However, this approach encounters a significant challenge: it fails when the number of tokens in a sequence is less than k, the predetermined number of sampled tokens. This limitation can cause the algorithm to break, especially when working with shorter sequences or in scenarios where the token count fluctuates.
To address this issue, we devised a solution involving the creation of 2k learnable key and value vectors, each associated with learnable importance scores. These vectors are sampled alongside the input tokens during the attention process. These additional vectors function similarly to support vectors in traditional kernel methods, providing additional decision boundaries within the self-attention mechanism. The introduction of these learnable support vectors not only mitigates the issue of insufficient tokens but also offers several other benefits. Firstly, they expand the learning capacity of transformers, allowing them to capture more nuanced patterns and relationships within the data. Secondly, these vectors stabilize the training process by ensuring that the model consistently attends to a fixed number of tokens, thereby maintaining the robustness and reliability of the attention mechanism across varying input sizes.
Mixture of Attention Experts.
While global attention mechanisms, such as the one we propose, offer significant advantages in terms of parallelizability and runtime efficiency, they are not without limitations. One key drawback is their restricted local receptive field, meaning that each token can only attend to the same subset of selected tokens within each layer. This limitation arises from the inherent design of global attention, where every token in the sequence attends to the same sampled tokens across all attention heads. However, upon closer examination, we realized that each attention head operates independently of the others. This independence implies that the outputs of the self-attention mechanism remain unchanged regardless of the token permutations applied within each attention head.
To overcome this limitation, we drew inspiration from the Mixture of Experts (MoE) literature, particularly the concept of expert-choosing-token (Zhou et al., 2022). Our method introduces a novel approach where each attention head selects a different subset of sampled tokens to attend to. This process can be executed in parallel, as the results of the h-sampling without replacement are independent of one another. By enabling each attention head to focus on a distinct subset of tokens, our transformer architecture effectively increases the local receptive field by a factor of h, where h represents the number of attention heads. This means that with 4 attention heads and a 25% sampling rate, our model can approximate a full self-attention pattern when necessary, significantly enhancing its ability to capture both global and local dependencies within the data.
Putting It All Together.
Our transformer architecture integrates several cutting-edge techniques to maximize performance and efficiency. We adopt the Pre-RMSNorm formulation (Chen et al., 2018; Zhang and Sennrich, 2019), which has been proven to be both faster and more conducive to a higher convergence rate during training. This formulation normalizes the input before applying the self-attention mechanism, leading to improved stability and efficiency. For the residual branch of our model, we employ a zero-initialized learnable scalar multiplier (Bachlechner et al., 2021), a technique known to enhance the training dynamics of deep networks. However, we introduce a slight modification: the learnable scalar multiplier is shared across all layers of the network, promoting consistency and reducing the number of parameters that need to be learned. For importance score computation, a multi-layer perceptron transforming tokens’ latent is sufficient because the latents are complexly transformed by multiple transformer layers before.
Modifications to integrate our sampling mechanism is described in earlier text. This integration allows our model to benefit from the efficiency and enhanced learning capacity provided by the differentiable top-k sampling and learnable support vector paddings. Finally, to align our model with the latest advancements in language technology, we incorporate flash attention (Dao et al., 2022; Dao, 2024). Flash attention is a highly optimized variant of self-attention that offers higher throughput and a lower memory footprint, making it ideal for large-scale models like ours. By combining these state-of-the-art techniques, our transformer architecture is not only capable of achieving top-tier performance across a variety of tasks but also does so with greater efficiency, ensuring it can be deployed effectively in real-world applications where computational resources and time are critical factors.
4.4. Complexity Analysis
@@
Consider the input token $\mathbb{X}\in\mathbb{R}^{n\times d}$, query and key weights $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$, and $k$ sampled tokens, the computational cost for each 1-head transformer layer is as follows:
  • Linear transformations for $\mathbb{Q},\mathbb{K}$ and $\mathbb{V}$: $O(nd^{2})$,
  • Importance score of tokens for each attention head: $O(nd)$,
  • Selecting top-$k$ importance score: $O(n+k)$ forward pass, $O(n+k^{2})$ backward pass with the formulation in Section 4.2,
  • Attention from between $n$ queries and top-$k$ most important keys: $O(nkd_{h})$,
@@
Given these complexities and considering the long-ranged data structures which can implies $n\gg d$, the asymptotic computational complexity of one single SAMSA layer is $O(n)$. The asymptotic space complexity of one single SAMSA layer is also $O(n)$.
5. Experiments
Table 1. Performance/Runtime of SAMSA compared to the full self-attention (Vaswani et al., 2017) with different sampling sizes. Green and Red indicate better and worse performances/runtimes respectively.
SAMSA Full Attention
Sampling Size 128 128 256 256 512 512 N/A
@@ 5. Experiments
Sequence ListOPS 41.53 41.13 41.89 40.57 41.33 40.88 37.11
Pathfinder 81.97 71.83 80.55 69.71 79.74 72.28 71.83
Image 48.64 48.24 48.73 47.38 48.24 46.75 42.94
Text 64.74 65.42 65.53 65.42 65.38 65.42 65.21
Retrieval - 82.05 - 81.89 - 81.81 79.14
Graph Pept-func 72.01 72.01 71.42 72.21 - 72.45 63.26
Pept-struct 0.2519 0.2458 - 0.2515 - 0.2468 0.2529
Point Cloud ModelNet40 91.11 90.91 91.72 90.95 91.35 91.39 91.07
ShapeNetPart - 85.74 - 85.87 - 85.62 85.1
Inference Speed (1024 tokens) 1.63×1.63$\times$1.63 × 2.17×2.17$\times$2.17 × 1.44×1.$4$ 1.62×1.62$\times$1.62 × 1.08×1.08$\times$1.08 × 1.44×1.$4$ 1$\times$
5.1. Experimental Design
Our SAMSA models utilize either hard or soft sampling techniques, determined by whether the reparameterization trick is employed or not. Soft sampling foregoes the reparameterization trick, instead operating with a very small Gumbel temperature (0.1), denoted as $g=g_{\text{soft}}$. This approach allows for a smoother, more probabilistic selection of tokens during training. Conversely, hard sampling employs the reparameterization trick with a standard Gumbel temperature (1.0), resulting in a more deterministic selection process. To rigorously evaluate the effectiveness of our approach, both in terms of performance and efficiency, we trained a total of 45 models. These models were benchmarked against full self-attention (specifically, Flash Attention v2 (Dao, 2024)) as well as other specialized state-of-the-art (SOTA) models across those three distinct data modalities:
  • Sequence (multilinear)
    • Long Range Arena (Tay et al., 2021)
  • Point Cloud (non-multilinear)
    • ModelNet40 (Wu et al., 2015)
    • ShapeNetPart (Chang et al., 2015)
  • Graph (non-multilinear)
    • Long Range Graph Benchmark (Peptides) (Dwivedi et al., 2022; Singh et al., 2015)
The selection of these datasets was intentional and rooted in a clear rationale. Our model, being a global attention model, was evaluated on datasets that highlight both the strengths and weaknesses typically associated with global attention mechanisms. Global attention models, while powerful in capturing long-range dependencies, are often criticized for their weakness in handling local structures. Several of the datasets we selected, such as 3 LRA datasets (Nangia and Bowman, 2018; Linsley et al., 2018; Krizhevsky and Hinton, 2009) and 2 LRGB datasets (Singh et al., 2015), are known to exhibit strong locality characteristics, making them ideal for testing this potential weakness. On the other hand, global attention models excel in tasks that require long-range modeling, a strength we tested using datasets that demand a global perspective, including 3 LRA datasets (Linsley et al., 2018; Maas et al., 2011; Radev et al., 2009), and 2 Point Cloud datasets (Wu et al., 2015; Chang et al., 2015).
Furthermore, the chosen data modalities—sequence, point cloud, and graph—are fundamentally different from each other, with minimal shared inductive biases. This diversity ensures that our model’s generalizability is rigorously tested, as it cannot rely on the exploitation of inherent structures unique to any one modality.
To ensure the validity and fairness of our comparisons with baseline models, we took several precautions to minimize bias. We avoided the use of data augmentation, which could introduce variability, and maintained nearly identical hyperparameters across different tasks. This consistent approach implies that we engaged in minimal hyperparameter tuning, allowing for a more accurate assessment of the SAMSA model’s inherent capabilities and its performance relative to other models under similar conditions. By controlling these variables, we aimed to present a clear and unbiased evaluation of SAMSA’s effectiveness across diverse data modalities and against a range of competitive benchmarks.
@@ 5.1. Experimental Design
Table 2. (Long Range Arena) Accuracy on the full suite of long range arena (LRA) tasks.
Method ListOps Text Retrieval Image Pathfinder
BigBird (Zaheer et al., 2020) 36.05 64.02 59.29 40.83 74.87
Reformer (Kitaev et al., 2020) 37.27 56.10 53.40 38.07 68.50
Performer (Choromanski et al., 2021) 18.01 65.40 53.82 42.77 77.05
Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34
Luna-256 (Ma et al., 2021) 37.98 65.78 79.56 47.86 78.55
SAMSA-hard (Ours) 41.12 65.42 81.89 48.24 72.27
SAMSA-soft (Ours) 41.88 65.53 82.05 48.73 81.97
5.2. Sampling vs. Full-Self Attention
We evaluated the performance of our SAMSA model across different sampling rates, with detailed results presented in Table 1. The experiments were conducted on nine different datasets using six distinct SAMSA model configurations. The results consistently demonstrate that SAMSA not only outperforms the full self-attention mechanism in most cases but also exhibits superior efficiency in both training and inference phases. This performance boost is particularly notable given the reduced computational overhead associated with SAMSA.
@@
Table 3. Experimental results on Peptides func and Peptides struct. We report the means and standard deviations of 4 random seeds.
@@
GatedGCN (Bresson and Laurent, 2017) 0.3420 ± 0.0013 0.5864 ± 0.0077
GatedGCN + RWSE (Bresson and Laurent, 2017) 0.3357 ± 0.0006 0.6069 ± 0.0035

### v2

@@ 4.2 Parallelizable Differentiable Top-k Sampling
We define $z$ as the importance score of vectors, representing the vectors’ likeliness of being picked. The importance scores of vectors are computed via a neural network: $\operatorname{ImportanceScore}(\mathbf{X}):\mathbb{R}^{n\times m}\rightarrow\mathbb{R}^{n}$:
$z\triangleq\operatorname{ImportanceScore}(\mathbb{X}).$
This simple method conditions the importance score of each token on the entire set of tokens, allowing patterns with complex dependencies to emerge. If we allow the set of sampled vectors to contain duplicates, Gumbel-Top-k algorithm can be parallelized simply by computing everything concurrently as follows:
$\operatorname{Sample}_{1}(z,\mathbf{X},k)$ $\triangleq\operatorname{Concatenate}_{i=1}^{k}(\operatorname{ST-GumbelSoftmax}(z)^{\top}\mathbf{X}).$
@@
The weakness of sampling with replacement is that it creates duplicates. These duplicates waste computational resources on the same pieces of information, which harms the expressivity of transformers relying on sampling. This proposed method of without-replacement sampling relies on a novel proxy soft gradient estimator we devised. To begin with, we started with a hypothetical inefficient but parallelizable method: select one set of sampled vectors from sets of all sets of sampled vectors using Gumbel-Softmax (this is the definition in Equation 3). This is equivalent to a continuous relaxation of the discrete brute-force combinatorial optimization algorithm. By defining the likeliness of choosing one set (i.e. the importance score of sets of vectors) as the sum of the importance scores of its elements, we have:
$z^{*}$ $\triangleq\operatorname{Concatenate}_{\mathbf{S}}^{C(\mathbb{N}^{\leq n},k)}(\sum_{i}^{\mathbf{S}}z_{i}),$
$\mathbf{X}^{*}$ $\triangleq\operatorname{Concatenate}_{\mathbf{S}}^{C(\mathbb{N}^{\leq n},k)}(\operatorname{vec}(\mathbf{X}_{\mathbf{S}})^{\top}),$
$\operatorname{Sample}_{2}(z,\mathbf{X},k)$ $\triangleq\operatorname{ST-GumbelSoftmax}(z^{*})^{\top}\mathbf{X}^{*},$
where $C(N,k)$ is the set of all k-combinations of set $N$, $\mathbb{N}^{\leq n}$ is the set of all natural numbers less than or equal to $n$, $\operatorname{vec}$ is the matrix vectorization function obtained by stacking the columns of the input matrix on top of one another, $\mathbf{X}_{\mathbf{S}}$ is the matrix of the row vectors with indices in $\mathbf{S}$ of $\mathbf{X}$. This definition of the importance scores of sets has a clear advantage: not needing to enumerate over the entire $C(n,k)$ at inference. Since using all possible combinations to generate one gradient vector for a single optimization step is overkill, an immediate improvement is local search - a sequential algorithm. Intuitively, this allows us to put the optimization steps of local search and gradient descent in the same for loop. For local search, we define the current solution as the set with the highest importance score. The locality of an index set $S$ is then described as the collection of index sets, each differing from $S$ by exactly one element:
$\operatorname{Current}$ $\triangleq\operatorname{argmax}(z^{*})=\operatorname{ArgTopK}(z)$
$\operatorname{Local}(\mathbf{S})$$\triangleq\{x:x\in C(\mathbf{X},k)\text{ and }\left|x-\mathbf{S}\right|=1\}$
Using local search as the base discrete combinatorial optimization algorithm, we compare the current solution against one of its localities, using Gumbel-Softmax reparameterization. The average result from comparing $k(n-k)$ pairs of solutions is given as follows:
$i$ $\triangleq\operatorname{ArgTopK}(z,k),$
$j$ $\triangleq\operatorname{ArgTopK}(-z,n-k),$
$p$ $\triangleq J_{k,n-k},$
$\partial p/\partial z$ $\triangleq\partial\left(\operatorname{ST-GumbelSigmoid}\left(z_{i}\mathbb{1}_{n-k}^{\top}-\mathbb{1}_{k}z_{j}^{\top}\right)\right)/\partial z,$
$\operatorname{Sample}(z,\mathbf{X},k)[m]$ $\triangleq\sum_{v=1}^{n-k}\frac{\mathbb{X}[i_{m}]*p[m,v]+\mathbb{X}[j_{v}]*(1-p[m,v])}{n-k},$
where $J_{m,n}$ is all-ones matrix with $m$ rows and $n$ columns, $\mathbb{1}_{k}$ are the all-one vectors of length $k$. The Gumbel Sigmoid function is the Gumbel Softmax when the number of categories is equal to two (details in Appendix D), similar to the relationship between Sigmoid and Softmax functions. To further cut down the optimization cost, we select $i$ as top-k most important and $j$ as top-k second most important index vectors.
4.3 Graph Modeling
The complexity of the graph data structure is needed for the claim of working on many data modalities of ours. Unfortunately, existing literature deemed self-attention alone insufficient for graph and they created very complicated methods: e.g. combination between Message Passing Neural Network (MPNN) and self-attention layers (Rampášek et al., 2022; Ying et al., 2021; Yun et al., 2019). This makes it hard to adapt to other data modalities. We identify the reason transformers need so many modifications just to be barely working on graphs is the inability to process both the nodes and edges information simultaneously. Therefore, we propose Graph-Bridge, a simpler approach treating both nodes and edges as tokens, which can be applied to any self-attention network without complicated constructions.
Formulation.
Given a graph $G=(V,E)$ where $V$ represents the set of nodes and $E$ represents the set of edges, the objective is to learn a mapping $\mathcal{F}:G\rightarrow\mathbb{R}^{d}$ that captures the structural properties of $G$ in a $d$-dimensional latent space. Each node $v_{i}\in V$ is associated with a feature vector $\mathbf{x}_{i}\in\mathbb{R}^{f}$, and each edge $e_{ij}=(v_{i},v_{j})\in E$ is associated with a feature vector $\mathbf{e}_{ij}\in\mathbb{R}^{g}$.
Positional Encoding.
To integrate node and edge information, we introduce a novel graph positional encoding (PE) mechanism. Let $\mathbf{p}_{i}\in\mathbb{R}^{d}$ denote the positional encoding of node $v_{i}$. We define $\mathbf{p}_{i}$ as a random vector sampled from a normal distribution:
$\mathbf{p}_{i}\sim\mathcal{N}(0,\operatorname{diag}(\sigma^{2}))+\phi(x_{i}),$
where $\sigma^{2}$ is a learnable variance $d$-dimensional vector and $\phi:\mathbb{R}^{f}\rightarrow\mathbb{R}^{d}$ is a multi-layer perceptrons (MLPs) to inject node feature into positional encodings. For each edge $e_{ij}$, its positional encoding $\mathbf{p}_{ij}$ is computed as:
$\mathbf{p}_{ij}=\mathbf{p}_{i}-\mathbf{p}_{j}.$
While summation is usually used to combine positional encodings, in this case, we use subtraction for edge encoding to preserve the directional information of graph edges.
Node and Edge Transformation.
To map node and edge representations into a common latent space, we use two separate MLPs, $\phi_{V}:\mathbb{R}^{d}\times\mathbb{R}^{f}\rightarrow\mathbb{R}^{d}$ and $\phi_{E}:\mathbb{R}^{d}\times\mathbb{R}^{g}\rightarrow\mathbb{R}^{d}$, defined as:
$\mathbf{h}_{i}$ $=\phi_{V}(\mathbf{p}_{i},\mathbf{x}_{i}),$
$\mathbf{h}_{ij}$ $=\phi_{E}(\mathbf{p}_{ij},\mathbf{e}_{ij}),$
where $\mathbf{h}_{i}$ and $\mathbf{h}_{ij}$ are the transformed representations of node $v_{i}$ and edge $e_{ij}$, respectively. These features are then processed by subsequent transformer layers, allowing complex relationships between nodes and edges to be recognized.
Compare to other Graph-PE.
This method is simple and efficient to compute: requiring only random feature generation, gathering operation on the randomly generated node PE for edges, and MLP features. This is different to existing graph positional encodings like Random-Walk PE (Ma et al., 2023a) (hard to parallelize) and Laplacian PE (Rampášek et al., 2022) (inherently computationally expensive). While being different, how this method works is surprisingly similar to random-walk PE: each self-attention propagates signals through edge tokens, i.e. we have merged the sequential process of random-walk and multi-layer neural network altogether.
4.4 Integration to Self-Attention Mechanism
While a straightforward integration of our sampling method into the self-attention mechanism is functional, we can achieve significant improvements. The following sections outline our enhancements.
Mixture of Attention Experts.
While global attention mechanisms are highly parallelizable and efficient, they limit each token’s local receptive field since all tokens attend to the same set of selected tokens. Each attention head operates independently, making self-attention outputs invariant to token permutations within each head. Inspired by MoE (Mixture of Experts) (Zhou et al., 2022), we propose a method where each attention head selects a distinct subset of sampled tokens. This process is parallelizable since the $h$-sampling processes are independent. Formally, the multi-head importance score $\mathbf{Z}$ is computed by:
$\mathbf{Z}\triangleq\operatorname{MultiHeadImportanceScore}(\mathbb{X}).$
And the multi-head sampling function $\operatorname{MultiHeadSample}$ as follows:
$\mathbf{\tilde{P}}$ $\triangleq\operatorname{MultiHeadSample}(\mathbf{Z},\mathbf{P},k)$
$=\operatorname{Stack}_{i=1}^{h}(\operatorname{Sample}(\mathbf{Z}_{\dots,i},\mathbf{P},k)),$
where $\mathbf{Z}_{\dots,i}$ is the $i^{\text{th}}$ column vector of $\mathbf{Z}$ matrix. By incorporating this mixture of attention experts, our sampling transformer significantly enhances its local receptive field by a factor of $h$. For instance, a transformer with four attention heads and a 25% sampling rate can learn a full self-attention pattern if necessary.
@@ 4.4 Integration to Self-Attention Mechanism
Simply attending all query vectors to a set of sampled key and value vectors has a limitation: when the number of tokens is less than k (the number of sampled tokens), our algorithm fails. To address this issue, we introduce 2k learnable key and value vectors, each associated with learnable importance scores, which are sampled concurrently with input tokens. The process can be described mathematically as follows:
$\mathbf{P}$ $\leftarrow\operatorname{Concatenate}(\mathbf{P},\mathbf{P_{\text{supp}}}),$
$\mathbf{Z}$ $\leftarrow\operatorname{Concatenate}(\mathbf{Z},\mathbf{Z}_{\text{supp}})$
where $\mathbf{P}:\mathbb{R}^{n\times 2d}$ is the key-value matrix, $\mathbf{Z}:\mathbb{R}^{n\times h}$ is the multi-head importance score matrix, $\mathbf{P_{\text{supp}}:\mathbb{R}^{2k\times 2d}}$ is the learnable support vector matrix, and $\mathbf{Z_{\text{supp}}:\mathbb{R}^{2n\times h}}$ is the learnable multi-head importance score matrix. These vectors and learned importance scores function is much similar to support vectors and Lagrange multipliers in traditional kernel methods, effectively introducing additional decision boundaries within the self-attention mechanism. Consequently, this not only enhances the learning capacity of transformers but also ensures stable training by consistently attending to a fixed number of tokens.
Architecture Details.
Our transformer architecture employs the Pre-RMSNorm formulation (Chen et al., 2018; Zhang & Sennrich, 2019), which has been demonstrated to accelerate convergence rates. For the residual branch, we utilize a zero-initialized learnable scalar multiplier (Bachlechner et al., 2021), with a modification: the learnable scalar multiplier is shared across all network layers. The self-attention modules in our architecture are adapted to incorporate our sampling mechanism, as discussed in detail. The importance score neural network for each self-attention module is implemented as a simple multi-layer perceptron, as the latent representations produced by the preceding transformer stack already exhibit sufficient complexity and expressiveness. Finally, to align with recent advancements in language modeling, we integrate flash attention (Dao et al., 2022; Dao, 2024) into our architecture, which offers higher throughput and reduced memory usage. The full algorithm is summarized in Algorithm 1, which is in the overview section.
4.5 Complexity Analysis
@@
For an input token $\mathbb{X}\in\mathbb{R}^{n\times d}$ with query and key weights $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{a}}$ and $k$ sampled tokens, the computational costs per 1-head transformer layer are:
  • Linear transformations for $\mathbb{Q},\mathbb{K},\mathbb{V}$: $O(nd^{2})$,
  • Calculating token importance scores: $O(nd)$,
  • Selecting top-k tokens: $O(n+k)$ forward, $O(n+k^{2})$ backward,
  • Attention between $n$ queries and top-k keys: $O(nkd_{h})$,
@@
For large number of tokens ($n\gg d$), the asymptotic time and space complexity of SAMSA layer is $O(n)$.
5 Experiments
5.1 Experiment Design
Our SAMSA models are either using hard/soft sampling indicating whether the straight through estimator is used or not. We have trained 45 models to measure the effectiveness of our method, both the performance and efficiency-wise of our model, against full self-attention (Flash Attention v2 (Dao, 2024)) and other specialized SOTA models on three data modalities (Sequence, Point Cloud, and Graph) on 4 datasets (Long Range Arena (all five), ModelNet40, ShapeNetPart, and Long Range Graph Benchmark (only Peptides)). To minimize any possible bias when comparing between baselines, we use no data augmentation and identical hyperparameters across different tasks. This also implies we conducted little hyperparameter tuning and mostly focused on the architectural design.
5.2 Performance-Efficiency Trade-offs
Table 1: Performance/Runtime of the full self-attention (Vaswani et al., 2017) compared to SAMSA with three different sampling sizes and two sampling formulations. Green and Red indicate SAMSA having better or worse statistics compared to full self-attention, respectively. The results of full self-attention in LRA are obtained from (Ma et al., 2023b). It has better performance than what reported in the dataset paper (Tay et al., 2021). The results of full self-attention in LRGB are obtained from the dataset paper (Dwivedi et al., 2022). The results of full self-attention on Shape Classification on ModelNet40 is run by ourselves, with the exact same hyperparameters as other SAMSA models. The results of PartSegmentation on ShapeNetPart is obtained from PointBERT (Yu et al., 2022).
SAMSA Full Self-Attention
Exp Settings Sampling Size 128 256 512 N/A
@@ 5.2 Performance-Efficiency Trade-offs
Sequence ListOPS L=2048 Acc 41.53 41.13 41.89 40.57 41.33 40.88 37.11
Pathfinder L=1024 Acc 81.97 71.83 80.55 69.71 79.74 72.28 71.83
Image L=1024 Acc 48.64 48.24 48.73 47.38 48.24 46.75 42.94
Text L=2048 Acc 64.74 65.42 65.53 65.42 65.38 65.42 65.21
Retrieval L=4096 Acc - 82.05 - 81.89 - 81.81 79.14
Graph Pept-func L=450 AP 72.01 72.01 71.42 72.21 - 72.45 63.26
Pept-struct L=450 MAE 0.2519 0.2458 - 0.2515 - 0.2468 0.2529
Point Cloud ModelNet40 L=1024 Acc 91.11 90.91 91.72 90.95 91.35 91.39 91.07
ShapeNetPart L=2500 IoU - 85.74 - 85.87 - 85.62 85.1
Inference Speed (1024 tokens) $\mathbf{1.63\times}$ $\mathbf{2.17\times}$ $\mathbf{1.44\times}$ $\mathbf{1.62\times}$ $\mathbf{1.08\times}$ $\mathbf{1.44\times}$ 1$\times$
We evaluated the performance of our SAMSA model on nine different datasets using six distinct SAMSA model configurations. The detailed results is presented in Table 1.
Trade-offs against Full Self-Attention.
At a 50% sampling rate, SAMSA is significantly faster and cheaper to compute than full self-attention, particularly due to the hard formulation, which avoids the need for gradient estimation during inference. Performance-wise, SAMSA consistently outperforms full self-attention, even at very low sampling rates (e.g., Retrieval: 128 / 4k). This superior performance is likely due to the regularization effect introduced by the sampling operator. Early in training, the random selection of tokens acts like DropAttention regularization (Zehui et al., 2019), preventing overfitting by ensuring attention is not overly concentrated on specific tokens. As training progresses, the sampler becomes more effective at identifying important tokens, reducing the regularization effect and allowing for more targeted attention. This dynamic aligns with known phenomena where early-stage regularization accelerates convergence, leading to improved overall performance (Liu et al., 2023b).
Trade-offs between Different Sampling Rates.
Efficiency-wise, Table 1 shows that SAMSA’s inference speed decreases with lower sampling rates and can be twice as fast as full self-attention on an A100. Performance-wise, the table also shows that a sampling rate of 256 tokens per head outperforms both full self-attention and higher sampling rates, suggesting that not all tokens equally impact attention effectiveness. This "sampling nonlinearity" may help filter out redundant tokens. Additionally, SAMSA’s performance on easy tasks like LRA-Text and LRA-Image decreases as sampling rates increase, indicating a potentially stronger regularization effect of lower sampling rates.
Trade-offs between Hard and Soft Sampling.
Hard-SAMSA is much faster than soft-SAMSA, with a 33% speed advantage at a 50% sampling rate. However, hard-SAMSA underperforms in certain tasks, likely due to discrepancies between the forward and backward passes in Gumbel-Softmax (Jang et al., 2017). Notably, hard-SAMSA’s learning curves in the Pathfinder task are volatile, reflecting the adverse effects of the straight-through estimator. Interestingly, earlier navigation through "breakpoints" leads to worse performance, possibly due to flat local minima, as shown by the stair-like learning curves (figure in Appendix E). However, it should be noted that hard-SAMSA performance is not inferior to soft-SAMSA with a counter-example: superior performance on graph data modality.
5.3 Sampling vs. Specialized Models
Aside from comparing against full self-attention, we also compare our model against the SOTA-specialized model on each of our experimented data modalities (sequences, graph, and point cloud).
Sequences.
@@
Table 2: (Long Range Arena) Accuracy on the full suite of long range arena (LRA) tasks. Bold indicates best sequence-models while underline indicates best data modality agnostic models.
Method ListOps Text Retrieval Image Pathfinder Modality
BigBird (Zaheer et al., 2020) 36.05 64.02 59.29 40.83 74.87 Sequence
Reformer (Kitaev et al., 2020) 37.27 56.10 53.40 38.07 68.50 Any
Performer (Choromanski et al., 2021) 18.01 65.40 53.82 42.77 77.05 Any
Linformer (Wang et al., 2020) 35.70 53.94 52.27 38.56 76.34 Any
Luna-256 (Ma et al., 2021) 37.98 65.78 79.56 47.86 78.55 Any
Transformer (Vaswani et al., 2017) 37.11 65.21 79.14 42.44 71.83 Any
S4 (Gu et al., 2022) 88.65 76.02 87.09 86.09 86.05 Sequence
MEGA (Ma et al., 2023b) 63.14 90.43 91.25 90.44 96.01 Sequence
hard-SAMSA (Ours) 41.12 65.42 81.89 48.24 72.27 Any
soft-SAMSA (Ours) 41.88 65.53 82.05 48.73 81.97 Any
In Table 2, our model has competitive results against many other efficient transformers and the full attention transformer. From the table, it can be seen that MEGA (Ma et al., 2023b) and S4 (Gu et al., 2022) have the highest performance. This is true and expected because recurrent and locality via EMA are very strong inductive bias, making the model only work for sequences. Therefore, it is natural that our method, which is designed for multiple-data modalities, cannot outperform methods like those. Against weaker to no inductive bias like the rest, we achieved better performance on every task except text classification (losing Luna-256 by 0.25%).
Graph.
@@
Table 3: Experimental results on Peptides func and Peptides struct. The results of methods other than ours are taken from their respective works and LRGB (Dwivedi et al., 2022). We report the means and standard deviations of 4 random seeds.
@@
GatedGCN (Bresson & Laurent, 2017) 0.3420 ± 0.0013 0.5864 ± 0.0077
GatedGCN + RWSE (Bresson & Laurent, 2017) 0.3357 ± 0.0006 0.6069 ± 0.0035